# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ 1. New avenue to STF or procedural issues
    The adoption of a BP creates a new gateway for parties to bring their cases to the STF. Since the existing legislation restricts access to the STF, these cases might otherwise have received their final decision from a lower court. Conversely, losing litigants have an incentive to take advantage of a new opportunity to reach STF in an attempt to change the lower court’s decision. This also creates procedural disputes regarding the proper use of specific classes of appeals and what cases fall under the new BP.
@@ 2. Resistance by a group of litigants or regional specificity
    A group of recurrent litigants or judges deviates from the pacified interpretation of the BP. Because of the current organization of the Brazilian Judiciary, it is not uncommon for regions and states to develop diverging positions on specific legal issues. When a new BP conflicts with such a position, many cases that reach STF might stem from this disagreement, following an attempt from specific groups to influence the interpretation of the BP.
@@ 4. New theme
    The BP introduced a new theme, resulting in a variety of cases discussing this previously unmentioned topic in STF. It should be noted that, in every legal system, judicial decisions entail a creative element, particularly heightened when a court possesses the authority to establish general norms through precedents. Thus, a BP might also be wielded as an instrument of judicial activism, whenever STF encounters a situation where current social practices are deemed contradictory to the Constitution.
@@ 4.2 BP 11
During the trial of HC (Habeas Corpus) 91.952555HC 91.952 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=570157 (08/07/2008), the defense requested the removal of handcuffs from the defendant, due to concerns about the negative perception that the sight of handcuffs could convey to the jury. This case triggered a highly publicized debate, which then shifted to discussing the use of handcuffs for public exposure, and more generally the sensationalization of criminal prosecution in the media. Based on the presumption of innocence, individual freedom, and the dignity of the human person, the High Court voted a few days later on the following:
@@
It is worth mentioning certain controversies surrounding the precedent. First, the nature of the nullity (absolute or relative) is not explicitly specified. Currently, the Supreme Court understands that the nature of the nullity is relative, as it depends on a demonstration of concrete harm to the defendant. However, [53] note that lower courts across the country have been deciding similar cases under the assumption of absolute nullity. Besides, as pointed out by [54] and [17], a question remains regarding the exact modalities of justification of the fear of escape or danger evoked in the text.
@@
Mathematical results
We show in Fig. 3 the predictions of the five models on our whole collection of documents (Dataset #2), as well as a regex search for documents containing the words algemas or algemado (in English, “handcuffs” and “handcuffed”). As explained in Section 4.1, the thresholds of the models, trained on Dataset #1, are not adapted for Dataset #2: way too many documents are predicted. Therefore, we choose new thresholds with the following rule: the highest value such that at least 90% of the documents citing BP 11 are predicted. In the figure, the initial and the tuned prediction curves are visualized as dashed or solid curves, respectively.
Models’ predictions for BP 11 Refer to caption
Figure 3: Number of documents predicted by each model for BP 11 in Dataset #2, represented as a histogram (window length of six months) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and #2 (solid), as described in LABEL:tab:adjusted_probabilities. Two views are given, the second one zooming in on the ordinate axis.
As it turns out, there are almost no documents, before the publication of BP 11, containing the words of the regex query. This phenomenon is also illustrated by the models TF-IDF+SVM, TF-IDF+logistic, and TF-IDF+randomforest, which closely follow the groundtruth curve. As a matter of fact, among the 853 documents of our dataset containing the words algemas or algemado, respectively 422, 433, and 420 have been predicted by the model, accounting for $79.9\%$, $63\%$ and $99.1\%$ of their predictions. In these documents, these words are often used in conjunction with uso and nulidade (use and nullity), also present in the official wording of BP 11.
The other models, however, estimate thousands of documents. Only 3.6% of LSTM’s predictions contain the words algemas or algemado, and 1.4% for BERT. A manual inspection shows that these models tend to detect documents merely containing the words uso or nulidade, independently of the surrounding context. For instance, among the 148,923 documents containing the word uso, the models LSTM and BERT predict, respectively, 6903 and 13,866 documents, accounting for 51,6% and 36,1% of predicted documents. This suggests that these models are biased toward detecting this word, regardless of its role in the conjunction uso de algemas (use of handcuffs).
To analyze this phenomenon further, let us consider the TF-IDF’s most relevant words, obtained in LABEL:tab:importance_features: algemado, algemas, audiencia, fuga, justificada, nulidade, and uso (handcuffed, handcuffs, audience, runaway, justified, nullity, use). We give in Fig. 4 the correlations between these seven words, when restricting the dataset to the documents explicitly citing BP 11, as well as the documents estimated by the models. As explained in Section 4.1, a well-fitted model is expected to show similar correlations as the ground truth. The first image shows a high correlation between uso and algemas, indicating that they mainly belong to the set uso de algema. Besides, only the TF-IDF-based models reflect this behavior. This observation accounts for the fact that LSTM and BERT, by not taking into account the correlation between these words, show poor results when used on the larger dataset.
Figure 4: Correlations between the selected relevant words, when restricted to documents citing BP 11 (“Groundtruth”) or documents predicted by one of our models. Please note that, as discussed in Section 4.1, correlations are calculated by viewing the words as binary random variables, hence correlation values lie between 0 and 1.
Juridical discussion
@@
As a consequence of the analysis above, we take TF-IDF models to be the most reliable for the retrieval of documents similar to those citing BP 11. We shall see in the next sections that this conclusion holds for all the precedents considered. The case of BP 11, however, is unfortunate: no document is predicted by the models before the creation of the precedent, hence no interesting behavior can be observed in Fig. 3. This restricts the application of our methodology described in Section 4.1. Nevertheless, two juridical insights can be drawn. First, the regex search shows that the term “handcuffs” has come to be used in STF’s cases recently, thanks to the edition of BP 11. It should be noted that a regex search for uso de força (use of force), a term commonly used as a synonym of handcuffs, only finds 153 documents, 74.8% of which also cite the precedent. On top of that, the fact that the TF-IDF models do not detect any similar documents before 2008 suggests that BP 11 not only changed the way cases are formulated, but really introduced this matter in the high court. This observation would fall under the New theme hypothesis, evoked in Section 4.1 as a way of explaining the inefficiency of the precedent in reducing the number of cases.
@@
We see in this precedent that a BP does not solely serve to pacify jurisprudence, but it might also be used as a tool of judicial activism by STF, to resolve a situation deemed contrary to the Constitution. Among these situations is the widespread use of handcuffs, addressed in the precedent, which, despite being widely accepted by the public, the Justices deemed to be incorrect and to violate the basic rights of the defendants. In this case, the BP did not originate from a recurrent legal controversy, but from the realization, by the court, that common social practices were in sharp contradiction with the Law.
@@
Following our methodology, we represent in Fig. 5 the type of juridical processes used in conjunction with BP 11. As observed, most of these cases are reclamações (complaints/appeals, abbreviated Rcl), a specific process filed to contest a decision or address a grievance, typically related to matters falling within the jurisdiction of the STF. To understand the purpose of these claims, we read some of the 632 documents citing BP 11. It turns out that many decisions of the STF are merely “procedural”, in the sense that they discuss the unsuitability of the reclamação as the proper avenue for the case. All these cases are deferred, based on internal procedural regulation. This seems to indicate that the reclamação is used by lawyers as a substitute for an appeal, corresponding to our hypothesis New avenue to the court and procedural issues.
Additional metadata for documents predicted by TF-IDF+randomforest for BP 11 Refer to caption
@@
Figure 5: Type of legal process used in the documents predicted by TF-IDF+randomforest for BP 11 in Dataset #2, among ARE (Recurso Extraordinário com Agravo), Rcl (Reclamação) and HC (Habeas Corpus).
@@
Besides, we point out that, among the reclamações that actually address the question of the use of handcuffs — i.e., non-procedural cases — the majority found their use adequately justified. However, these decisions often refrain from delving into a detailed discussion of the justification. This issue is reinforced by the legal precedent of preclusão (preclusion), which states that if there were no objections at the time of handcuff use, and if an authority figure was present, then the reclamação cannot be used. The often sparse written justification of handcuffs use, as well as the use of preclusão, call for a clarification of BP 11’s statement, as already mentioned in legal literature [54, 17].
@@ 4.3 BP 14
Binding Precedent 14 derived directly from the Attorney Statute of 1994666Article XIV https://www.oab.org.br/Content/pdf/LegislacaoOab/Lei-8906-94-site.pdf, which predicts the defense attorney’s rights “to examine, in any institution responsible for conducting investigations, even without authorization, records of flagrant crimes and investigations of any kind, completed or ongoing, even if already concluded for the authority, being able to copy documents and take notes, in physical or digital format”.
This right, however, is not free from dispute, especially in situations when classified investigations take place. The habeas corpus 88.190777HC 88.190 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=382091 exemplifies these further complications: in 2005, the Brazilian newspaper O Globo brought light to the existence of an investigation aiming at one of the owners of an important exportation company, a fact that was also new to the investigation’s target, given its classified status. After consistent refusals of investigation disclosure, lawyers decided to file an habeas corpus to the Federal Court of the Second Region (Tribunal Regional Federal da 2ª Região, TRF-2) against the Prosecutor’s Office, asking for access to the investigation proceedings. After failing to obtain a favorable decision at TRF-2 and the Supreme Court of Justice [^8] Different from the STF, this institution is the highest court of appeal in cases where no direct constitutional rights seem to be trespassed. (Superior Tribunal de Justiça), they filed an habeas corpus to STF pleading the recognition of unconstitutionality of the denial of access to the ongoing investigation documents, based on the guarantees of the right of sustaining contradictory positions through the whole criminal litigation. After debates, the STF Justices decided to allow the responsible attorneys to access the classified materials regarding their clients.
@@
Mathematical results
A quick analysis of the predicted documents in the bottom part of Fig. 6 indicates that, among the five models considered, only TF-IDF+random forest achieved satisfactory results, by faithfully following the groundtruth curve. This is also the case for the regex search of the expression acesso aos elementos/autos/documentos (access to elements, records, or documents).
Models’ predictions for BP 14 Refer to caption
Figure 6: Number of documents predicted by each model for BP 14 in Dataset #2, represented as a histogram (window length of six months) and interpolated via quadratic spline. We give the predictions for the thresholds adapted to Dataset #1 (dashed) and #2 (solid), as described in LABEL:tab:adjusted_probabilities. Two views are given, the second one zooming in on the ordinate axis.
To investigate why the other models estimate many documents, we will not consider the correlations, as we did for BP 11 (in Fig. 4), but simply the frequencies of the most relevant words, represented in Fig. 7. Almost all groundtruth documents cite acesso aos autos (access to records). Similarly, 96.3% of TF-IDF+randomforest’s predictions mention autos, of which 81.8% also mention acesso. Besides, among TF-IDF+SVM and TF-IDF+logistic’s predictions, respectively, 93.5% and 93.6% mention autos, while only 44.6% and 41.8% of them conjointly contain acesso. This suggests an underfitting of both models.
On the other hand, LSTM and BERT predict a quantity of documents not even mentioning the words acesso and autos. As it turns out, the decisions detected before the creation of BP 14 seemed of little relevance regarding its topic. Most of them999such as MS 26.024 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=435164 and HC 83.703 https://portal.stf.jus.br/processos/detalhe.asp?incidente=2180019 are related to “investigations”, although not addressing directly the attorneys’ right of access to the investigation material. We consider these predictions to be unreliable.
Interestingly, both the predictions of TF-IDF+SVM and TF-IDF+logistic peaked in 2017. At first sight, a manual inspection of a sample of the detected documents, excluding elements of the groundtruth class, raises questions about the qualities of these predictions. Although certain decisions [^10] Such as Inq 4.260 https://portal.stf.jus.br/processos/downloadPeca.asp?id=313917544&ext=.pdf and RHC 133.298 https://portal.stf.jus.br/processos/downloadPeca.asp?id=315014507&ext=.pdf are on topics regarding investigations (specifically, on the initiation of an official investigation and the transcription of some telephonic tapping record), neither of them concerns topics related to access of investigation material by the investigated part, which is the subject of the BP. The addition of metadata, as in Fig. 8, shows that this peak mainly consists of cases from Distrito Federal that also mention the word sigilo (secrecy/confidentiality). This suggests that these cases are complaints filed by attorneys representing individuals investigated or condemned for white-collar crimes, most relevantly, by attorneys representing politicians or individuals of political interest. It is, therefore, unsurprising that this increase in usage of the BP after 2012 partly matches the interval when Operação Lava Jato took place, a joint cooperative investigation aiming at politicians accused of corruption-related crimes. We continue this discussion, from a legal point of view, in the next paragraph.
Additional metadata for documents predicted by TF-IDF+SVM for BP 14 Refer to caption
Figure 8: Top: State to which the cases predicted by TF-IDF+SVM for BP 14 belong, with label “UNKNOWN” when the metadata is unavailable. Bottom: Regex search for the presence of the words sigilo (secrecy/confidentiality) and inquérito (inquiry/investigation).
Juridical discussion
@@
First, we shall analyze the predictions of models TF-IDF+SVM and TF-IDF+logistic, which, although not following the ground truth curve, evoke a political event worth pointing out. A reading of the documents shows that the peak of Fig. 8 can be dated to April 4th, 2017, the day which Judge Edson Fachin decided on 255 cases to proceed with investigations of criminal cases of corruption against holders of federal political positions, such as deputies and senators, based on denunciation of participation in the Lava Jato scheme. These decisions were pushed by the public prosecutors’ office’s accusation, based on information retrieved from leniency deals involving those others investigated through a special criminal institution known as delação premiada (turns state’s evidence). However, in the law establishing this form of testimony (12,850/2013111111https://www.planalto.gov.br/ccivil_03/_ato2011-2014/2013/lei/l12850.htm), secrecy is imposed on the shared information and bargain terms, aiming to preserve both the collaborator’s and the investigation’s integrity. The cases of April 4th, on the other hand, had important public and media impact, together with the many distinct requests from the public prosecutors’ office involving the establishment and continuation of investigation on individuals, petitions for publicity of the content of delations, and the list of those investigated. Not only did Judge Fachin decide to make the information public, but justified so, in all 255 documents, using very similar language. It is likely that such a “copy-and-paste” extract of text, in which the decider argues that the publicity of the investigation information does not detriment “the defense’s right, after collecting the accusatory piece, and with the means and financial resources to the adversary, the possibility of appealing against a complaint” — which is the topic of BP 14, although the decision, as a whole, is only marginally related to this issue — caused the models to detect some of these instances as potential applications of the BP.
@@
We now turn to the predictions of TF-IDF+random forest, which can be considered accurate, in view of their agreement with the groundtruth curve in Fig. 6. We gather additional metadata in Fig. 9, showing that BP 14 was not enough to settle all debates about lawyers’ access to investigation material, as already predicted right after its publication by [52]. By manually inspecting the decisions citing BP 14, we identified several complaints in which the defense of individuals, although not directly investigated, requested access to classified investigation material, due to some potential relation of their clients with the investigated individuals. Although unaddressed by the statement of the BP, the occurrence of this sort of complaint is frequent enough that the STF itself considered it as a subtopic of the BP [^12] See https://portal.stf.jus.br/jurisprudencia/sumariosumulas.asp?base=26&sumula=1230., with decisions going in both directions. The fact that BP 14 is cited in many cases not directly related to its content fits the hypothesis New avenue to STF or procedural issues.
@@
Figure 9: First: Type of legal process used in the documents predicted by TF-IDF+randomforest for BP 14, among HC (Habeas Corpus), Inq (Inquérito) and Rcl (Reclamação). Second: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Third: Type of decision taken at trial: processual if deferred for procedural matters, or deferido (accepted) and indeferido (rejected). Fourth: Regex search for the presence of the words corrupção (corruption), lavagem (money laundering), sigilo (secrecy/confidentiality) and eleitoral (electoral).
@@
This showcases how the realities of the Judiciary Power may be less predictable than most juridical theories on BPs seem to suggest. Although BPs are created to address specific legal problems, their subsequent use, in actuality, might be greatly influenced by new dynamics and unforeseeable trends in Case Law. Thus, as a greater number of BPs are created over time, the complexity of the legal system tends to increase and more opportunities arise for external factors to influence the usage of this legal instrument.

### v3

@@ 1. New avenue to STF or procedural issues
    The adoption of a BP creates a new gateway for parties to bring their cases to the STF. Since the existing legislation restricts access to the STF, these cases might otherwise have received their final decision from a lower court. Conversely, losing litigants have an incentive to take advantage of a new opportunity to reach the STF in an attempt to change the lower court’s decision. This also creates procedural disputes regarding the proper use of specific classes of appeals and what cases fall under the new BP.
@@ 2. Resistance by a group of litigants or regional specificity
    A group of recurrent litigants or judges deviates from the pacified interpretation of the BP. Because of the current organization of the Brazilian Judiciary, it is not uncommon for regions and states to develop diverging positions on specific legal issues. When a new BP conflicts with such a position, many cases that reach the STF might stem from this disagreement, following an attempt from specific groups to influence the interpretation of the BP.
@@ 4. New theme
    The BP introduced a new theme, resulting in a variety of cases discussing this previously unmentioned topic in the STF. It should be noted that, in every legal system, judicial decisions entail a creative element, particularly heightened when a court possesses the authority to establish general norms through precedents. Thus, a BP might also be wielded as an instrument of judicial activism, whenever the STF encounters a situation where current social practices are deemed contradictory to the Constitution.
@@ 5.2 BP 11
During the trial of HC (Habeas Corpus) 91.952131313HC 91.952 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=570157 (08/07/2008), the defense requested the removal of handcuffs from the defendant, due to concerns about the negative perception that the sight of handcuffs could convey to the jury. This case triggered a highly publicized debate, which then shifted to discussing the use of handcuffs for public exposure, and more generally the sensationalization of criminal prosecution in the media. Based on the presumption of innocence, individual freedom, and the dignity of the human person, the Supreme Court voted a few days later on the following:
@@
It is worth mentioning certain controversies surrounding the precedent. First, the nature of the nullity (absolute or relative) is not explicitly specified. Currently, the Supreme Court understands that the nature of the nullity is relative, as it depends on a demonstration of concrete harm to the defendant. However, [54] note that lower courts across the country have been deciding similar cases under the assumption of absolute nullity. Besides, as pointed out by [56] and [18], a question remains regarding the exact modalities of justification of the “fear of escape or danger” evoked in the text.
@@
Discussion
@@
In order to analyze BP 11 quantitatively, we return to the prediction curves, already analyzed in Section 4.2 (see Fig. 7). As we have seen, except regex after 2016 and LSTM, all the models faithfully follow the citation curve of the precedent after its publication, which means they are reliable for the retrieval of documents similar to those citing the binding precedent. The case of BP 11, however, is unfortunate: no document is predicted by the models before its creation, hence no interesting behavior can be observed in Fig. 7. This restricts the application of our methodology described in Section 5.1. Nevertheless, two juridical insights can be drawn. First, the regex search shows that the term “handcuffs” has come to be used in STF’s cases recently, thanks to the edition of BP 11. It should be noted that a regex search for uso de força (use of force), a term commonly used as a synonym of handcuffs, only finds 121 documents, 87.7% of which also cite the precedent. On top of that, the fact that the TF-IDF models and Longformer did not detect any similar documents before 2008 suggests that BP 11 not only changed the way cases are formulated but really introduced this matter in the Supreme Court. As a matter of fact, when reduced to branches “Criminal Procedure Law” and “Criminal Law”, Dataset #2 contains 20.3% of its documents published before 2008, a non-negligible proportion. This observation falls under the New theme hypothesis, evoked in Section 5.1 as a way of explaining the inefficiency of the precedent in reducing the number of cases.
@@
We see in this precedent that a BP does not solely serve to pacify jurisprudence, but it might also be used as a tool of judicial activism by the STF, to resolve a situation deemed contrary to the Constitution. Among these situations is the widespread use of handcuffs, addressed in the precedent, which, despite being widely accepted by the public, the Justices deemed to be incorrect and to violate the basic rights of the defendants. In this case, BP 11 did not originate from a recurrent legal controversy, but from the realization, by the court, that common social practices were in sharp contradiction with the Law.
@@
To provide a more detailed picture of BP 11’s applications, and following our methodology, we represent in Fig. 17 the type of juridical processes used in conjunction with the precedent. As observed, most of these cases are reclamações (complaints/appeals, abbreviated Rcl), a specific process filed to contest a decision or address a grievance, typically related to matters falling within the jurisdiction of the STF. To understand the purpose of these claims, we read some of the 830 documents citing BP 11. It turns out that many decisions of the STF are merely “procedural”, in the sense that they discuss the unsuitability of the reclamação as the proper avenue for the case. All these cases are deferred, based on internal procedural regulation. This seems to indicate that the reclamação is used by lawyers as a substitute for an appeal, corresponding to our hypothesis New avenue to the court and procedural issues.
Additional metadata for documents predicted by TF-IDF+SVM for BP 11 Refer to caption
@@
Figure 17: Type of legal process used in the documents predicted by TF-IDF+SVM for BP 11 in Dataset #2, among Rcl (Reclamação), ARE (Recurso Extraordinário com Agravo), and HC (Habeas Corpus).
@@
Besides, we point out that, among the reclamações that actually address the question of the use of handcuffs — i.e., non-procedural cases — the majority found their use adequately justified. However, these decisions often refrain from delving into a detailed discussion of the justification. This issue is reinforced by the legal precedent of preclusão (preclusion), which states that if there were no objections at the time of handcuff use, and if an authority figure was present, then the reclamação cannot be used. The often sparse written justification of handcuffs use, as well as the application of preclusão, call for a clarification of BP 11’s statement, as already mentioned in legal literature [56, 18].
@@ 5.3 BP 14
Binding Precedent 14 derived directly from the Attorney Statute of 1994141414Attorney Stat., Article XIV https://www.oab.org.br/Content/pdf/LegislacaoOab/Lei-8906-94-site.pdf, which predicts the defense attorney’s rights “to examine, in any institution responsible for conducting investigations, even without authorization, records of flagrant crimes and investigations of any kind, completed or ongoing, even if already concluded for the authority, being able to copy documents and take notes, in physical or digital format”.
This right, however, is not free from dispute, especially in situations where classified investigations take place. The Habeas Corpus 88.190151515HC 88.190 https://redir.stf.jus.br/paginadorpub/paginador.jsp?docTP=AC&docID=382091 exemplifies these further complications: in 2005, the Brazilian newspaper O Globo brought light to the existence of an investigation aiming at one of the owners of an important exportation company, a fact that was also new to the investigation’s target, given its classified status. After consistent refusals of investigation disclosure, lawyers decided to file a Habeas Corpus to the Federal Court of the Second Region (Tribunal Regional Federal da 2ª Região, TRF-2) against the Prosecutor’s Office, asking for access to the investigation proceedings. After failing to obtain a favorable decision at TRF-2 and the Supreme Court of Justice [^16] Different from the STF, this institution is the highest court of appeal in cases where no direct constitutional rights seem to be trespassed. (Superior Tribunal de Justiça), they filed another Habeas Corpus to the STF pleading the recognition of unconstitutionality of the denial of access to the ongoing investigation documents, based on the guarantees of the right of sustaining contradictory positions through the whole criminal litigation. After debates, the STF Justices decided to allow the responsible attorneys to access the classified materials regarding their clients.
@@
Discussion
@@
Let us turn to the predictions curves for BP 14, already presented in Fig. 9. First, we shall analyze the predictions of the models TF-IDF+SVM, TF-IDF+logistic, LSTM, and Longformer, which, although not following the groundtruth curve, evoke a political event worth pointing out. A reading of the documents shows that their common peak, visualized in Fig. 11, can be dated to April 4th, 2017, the day which Justice Edson Fachin decided on 255 cases to proceed with investigations of criminal cases of corruption against holders of federal political positions, such as deputies and senators, based on denunciation of participation in the Lava Jato scheme. These decisions were pushed by the public prosecutors’ office’s accusation, based on information retrieved from leniency deals involving those others investigated through a special criminal institution known as delação premiada (turns state’s evidence). However, in the law establishing this form of testimony (12.850/2013171717Law 2.850/2013 https://www.planalto.gov.br/ccivil_03/_ato2011-2014/2013/lei/l12850.htm), secrecy is imposed on the shared information and bargain terms, aiming to preserve both the collaborator’s and the investigation’s integrity. The cases of April 4th, on the other hand, had important public and media impact, together with the many distinct requests from the Public Prosecutors’ Office involving the establishment and continuation of investigation on individuals, petitions for publicity of the content of delations, and the list of those investigated. Not only did Justice Fachin decide to make the information public, but justified so, in all 255 documents, using very similar language. It is likely that such a “copy-and-paste” extract of text, in which the decider argues that the publicity of the investigation information does not detriment “the defense’s right, after collecting the accusatory piece, and with the means and financial resources to the adversary, the possibility of appealing against a complaint” — which is the topic of BP 14, although the decision, as a whole, is only marginally related to this issue — caused the models to detect some of these instances as potential applications of the BP.
@@
We now turn to the predictions of regex and TF-IDF+random forest, which can be considered accurate, given their agreement with the groundtruth curve in Fig. 9. We gather additional metadata in Fig. 18, showing that BP 14 was not enough to settle all debates about lawyers’ access to investigation material, as already predicted right after its publication by [53]. By manually inspecting the decisions citing BP 14, we identified several complaints in which the defense of individuals, although not directly investigated, requested access to classified investigation material, due to some potential relation of their clients with the investigated individuals. Although unaddressed by the statement of the BP, the occurrence of this sort of complaint is frequent enough that the STF itself considered it as a subtopic of the binding precedent [^18] BP 14, see Jurisprudência selecionada (Selected case law) https://portal.stf.jus.br/jurisprudencia/sumariosumulas.asp?base=26&sumula=1230., with decisions going in both directions. The fact that BP 14 is cited in many cases not directly related to its content fits the hypothesis New avenue to STF or procedural issues.
@@
Figure 18: Top: Type of legal process used in the documents predicted by TF-IDF+random forest for BP 14, among Rcl (Reclamação), Pet (Petição), and HC (Habeas Corpus). Middle: State to which these cases belong, with label “UNKNOWN” when the metadata is unavailable. Bottom: Type of decision taken at trial: processual if deferred for procedural matters, or deferido (accepted) and indeferido (rejected). In all three graphs, the bins have a length of six months.
@@
This showcases how the realities of the Judiciary Power may be less predictable than most juridical theories on binding precedents seem to suggest. Although BPs are created to address specific legal problems, their subsequent use, in actuality, might be greatly influenced by new dynamics and unforeseeable trends in Case Law. Thus, as a greater number of BPs are created over time, the complexity of the legal system tends to increase and more opportunities arise for external factors to influence the usage of this legal instrument.